# PAtt-Lite: Lightweight Patch and Attention MobileNet for Challenging Facial Expression Recognition

### v1

@@ 4.4.3 Results on FER2013
The comparison of the proposed method with the state-of-the-art methods on FER2013 is depicted in Table VI. FER2013 has been a database that existing work struggled to perform well on until recently when FLEPNet[51] and NECM-PECM Ensemble[52] were proposed. Compared to the existing work, PAtt-Lite recorded a classification performance of 92.50%, which equates to an improvement of 4.5% over NECM-PECM Ensemble[52].
@@
4.4.4 Results on FERPlus
FERPlus has an additional Contempt expression than the other in-the-wild databases in this research. The classification performance of our proposed method is compared with the state-of-the-art methods in the last column of Table IV. The accuracy and parameter of the proposed PAtt-Lite are also compared with state-of-the-art methods in Fig. 8. Like RAF-DB, transformer-based methods generally have better results than CNN-based methods, except for CIAO[46], which reported the best accuracy for FERPlus with a CNN backbone.
The proposed PAtt-Lite has achieved slight improvement on FERPlus when compared to the state-of-the-art methods. Specifically, PAtt-Lite achieved a classification accuracy of 95.55%, which corresponds to a performance improvement of 1.05% over CIAO[46], while having 16 times lesser parameters than CIAO[46]. When compared to transformer-based methods, our method achieved a 3.93% improvement in classification accuracy over POSTER[28] with 65 times lesser parameters. For comparison with lighter state-of-the-art methods, PAtt-Lite also outperformed RAN[5] by 6.39% in terms of overall accuracy with 10 times lesser parameters.
It is visible that our proposed method is still struggling with the Contempt class upon inspecting the confusion matrix in Fig. 6c, with the Disgust class and the Fear class following a general trend from RAF-DB and FER2013. However, this performance drop is expected due to the severe class imbalance between these classes and the Neutral class. Similar performance drops on some of these negative expression classes are also reported in existing work[15, 29, 28]. Despite this, our proposed method still achieved significant improvements in all other expression classes, with around or higher than 95% performance across these 5 classes.
@@ 4.4.4 Results on FERPlus
4.4.5 Results on challenging subsets
TABLE VII: Comparison of the state-of-the-art results on challenging subsets for RAF-DB and FERPlus. The best result is highlighted in bold.
@@ 4.4.5 Results on challenging subsets
The proposed method has shown great results in classifying samples under challenging conditions. Table VII shows the difference in the performance of the proposed method compared to recent work on challenging subsets for RAF-DB and FERPlus. Specifically, PAtt-Lite achieved an improvement of more than 3% across all three challenging subsets for RAF-DB, outperforming Facial Chirality[45] which reported the best results on these subsets with 46.2M parameters. Meanwhile, to our best knowledge, our proposed method is the first to achieve more than 90% accuracy for all challenging subsets for FERPlus. A performance improvement of around 4% to 7% is achieved across the three subsets compared to SCAN-CCI[15], which previously reported the best performance with 70M parameters. From the comparison with the best results on these subsets, the proposed PAtt-Lite achieved state-of-the-art performance across all subsets with significantly lesser parameters. Meanwhile, when compared with the lighter methods such as RAN[5] and Imponderous Net[16] on subsets of RAF-DB, PAtt-Lite managed to outperform these methods by around 9%. The proposed method also managed to outperform the lighter methods by more than 9% on the first two subsets and by 7.75% on the Pose 45 subset of FERPlus.
5 Conclusion
@@
This work provides valuable insight into potential future directions by highlighting the advantages and possible improvements of our proposed method. One such direction is robustness improvement to PAtt-Lite, especially towards low-resource expressions, which can be done by further refining the patch extraction blocks. We also suggest that there could be an effort for the development of a FER database with more reliable annotations. Existing FER methods have advanced to a point where they can potentially discover patterns that are difficult for human annotators to detect. Therefore, a crowdsourced database that incorporates insights from state-of-the-art FER methods could be a valuable resource for further research in this field.

### v2

@@ IV-E3 Results on FER2013
The comparison of the proposed method with the state-of-the-art methods on FER2013 is depicted in Table X. FER2013 has been a database that existing work struggled to perform well on until recently when FLEPNet[55] and NECM-PECM Ensemble[56] were proposed. Compared to the existing work, PAtt-Lite recorded a classification performance of 92.50%, which equates to an improvement of 4.5% over NECM-PECM Ensemble[56].
@@
IV-E4 Results on FERPlus
FERPlus has an additional Contempt expression than the other in-the-wild databases in this research. The classification performance of our proposed method is compared with the state-of-the-art methods in the last column of Table VIII. The accuracy and parameter of the proposed PAtt-Lite are also compared with state-of-the-art methods in Fig. 10. Like RAF-DB, transformer-based methods generally have better results than CNN-based methods, except for CIAO[48], which reported the best accuracy for FERPlus with a CNN backbone.
The proposed PAtt-Lite has achieved slight improvement on FERPlus when compared to the state-of-the-art methods. Specifically, PAtt-Lite achieved a classification accuracy of 95.55%, which corresponds to a performance improvement of 1.05% over CIAO[48], while having 16 times lesser parameters than CIAO[48]. When compared to transformer-based methods, our method achieved a 3.93% improvement in classification accuracy over POSTER[25] with 65 times lesser parameters, and a 2.46% improvement over ARBEx[47], a modified POSTER++[27]. For comparison with lighter state-of-the-art methods, PAtt-Lite with significantly lesser parameters also outperformed RAN[5], EAC[43], and DDAMFN[46] by 6.39%, 5.91%, and 4.81%, respectively, in terms of overall accuracy.
It is visible that our proposed method is still struggling with the Contempt class upon inspecting the confusion matrix in Fig. 8c, with the Disgust class and the Fear class following a general trend from RAF-DB and FER2013. However, this performance drop is expected due to the severe class imbalance between these classes and the larger class. Similar performance drops on some of these negative expression classes are also reported in existing work[9, 26, 25]. Despite this, our proposed method still achieved significant improvements in all other expression classes, with around or higher than 95% performance across these 5 classes.
@@ IV-E4 Results on FERPlus
IV-E5 Results on challenging subsets
TABLE XI: Comparison of the state-of-the-art results on challenging subsets for RAF-DB and FERPlus. The best result is highlighted in bold.
@@ IV-E5 Results on challenging subsets
The proposed method has shown great results in classifying samples under challenging conditions. Table XI shows the difference in the performance of the proposed method compared to recent work on challenging subsets for RAF-DB and FERPlus. Specifically, PAtt-Lite achieved an improvement of more than 3% across all three challenging subsets for RAF-DB, outperforming Facial Chirality[45] which reported the best results on these subsets with 46.2M parameters. Meanwhile, to our best knowledge, our proposed method is the first to achieve more than 90% accuracy for all challenging subsets for FERPlus. A performance improvement of around 4% to 7% is achieved across the three subsets compared to SCAN-CCI[9], which previously reported the best performance with 70M parameters. From the comparison with the best results on these subsets, the proposed PAtt-Lite achieved state-of-the-art performance across all subsets with significantly lesser parameters. Meanwhile, when compared with the lighter methods such as RAN[5] and Imponderous Net[10] on subsets of RAF-DB, PAtt-Lite managed to outperform these methods by around 9%. The proposed method also managed to outperform the lighter methods by more than 9% on the first two subsets and by 7.75% on the Pose 45 subset of FERPlus.
V Conclusion
@@
This work provides valuable insight into potential future directions by highlighting the advantages and possible improvements of our proposed method. One such direction is robustness improvement to PAtt-Lite, especially towards low-resource expressions, which can be done by further refining the patch extraction blocks or experimenting with a feature extractor backbone that has more performance or is more robust. Besides, as highlighted in the previous section, while large FER databases are available, good annotations are required to further advance the field of FER. Hence, we also suggest that there could be an effort for the development of a FER database with more reliable annotations. Existing FER methods have advanced to a point where they can potentially discover patterns that are difficult for human annotators to detect. Therefore, a crowdsourced database that incorporates insights from state-of-the-art FER methods could be a valuable resource for further research in this field.